OpenStudy (anonymous):

3x^-2=2x^4 explain :)

OpenStudy (anonymous):

\[3x^{-2}=2x^4\] Right??

OpenStudy (anonymous):

\[x^{-a} = \frac{1}{x^a}\]

OpenStudy (anonymous):

@waterineyes yes thats right

OpenStudy (anonymous):

So : \[\frac{3}{x^{2}} = 2x^4\] Now, multiply by \(x^2\) both the sides and tell me what did you get??

OpenStudy (anonymous):

3=2x^4*x^2 ???

OpenStudy (anonymous):

not so sure..

OpenStudy (anonymous):

Yep but why are not sure??

OpenStudy (anonymous):

Now : \[x^a \cdot x^b = x^{a+b}\]

OpenStudy (anonymous):

That was extremely confusing.. what do those values represent ? sorry for the lack of ability to do this.. sort of a beginner on this math..